Consider the following hypothesis test. a. What is your conclusion if n 1-21, s 1 2-87, n 2 26, and s 2 2 4? Use α-05 and the p-value approach. Calculate the value of the test statistic (to 2 decimals). 2.05 3 The p-value is between .02 and.05 What is your conclusion? Cannot conclude that the two population variances are different b. Repeat the test using the critical value approach. What is the rejection rule for this hypothesis test? Round the critical values to 2 ecimal places. Reject Ho if F greater than 05 What is your conclusion? Cannot conclude that the two population variances are different
Answer #1

We are to conduct a 2 sample test of equality of variances.

Given, n1=21 s12=8.7,n2=26 ,s22=4

The test statistic F is given by ,

                                               F = (n1-1)s12 / (n2-1)s22

                                                                  = (21-1)*8.7/ (26-1)*4

                                                   = 174 / 100

                                                   = 1.74

The p-value is greater than 0.05.

We reject Ho if the p-value is less than \alpha.

The rest of the answers as marked by you are correct.